OpenStudy (anonymous):

Solve 5c - c + 10 = 34.

OpenStudy (studygurl14):

Add/subtract like terms: 5c - c + 10 = 34 4c = 34 - 10 4c = 24 c = ?

HanAkoSolo (jamierox4ev3r):

add the like terms here. If you do so, you should come to an answer :)

HanAkoSolo (jamierox4ev3r):

DO you understand how to go about doing this? @LazyDirectioner

OpenStudy (the_fizicx99):

5c - c + 10 = 34. Combine like terms 5c - c = 4c 4c + 10 = 34 -10 -10 4c = 24 /4 /4 c = ??

OpenStudy (anonymous):

@Jamierox4ev3r yes i got the answer its C=6

OpenStudy (the_fizicx99):

Check: 5(6) - 6 + 10 = 34. 30 - 6 + 10 24 + 10 = 34 34 = 34 :P

OpenStudy (anonymous):

thanks @tHe_FiZiCx99 @Jamierox4ev3r @StudyGurl14

OpenStudy (studygurl14):

yw